OpenStudy (babyslapmafro):

Please help me solve the following problem. Find the local linear approximation L to the specified function f at the designated point P(4,3). click to see function

OpenStudy (babyslapmafro):

\[f(x,y)=\frac{ 1 }{ \sqrt{x ^{2}+y ^{2}} }\]

OpenStudy (anonymous):

alright if you don't mind me giving a shot at it i have not gone into calcu. yet we are trying to find the tangent of the point (4,3) but with the given function f(x,y) we can find it tangent at x = a and i think that the equation for the tangent line is \[L(x) = f(a) + f \prime (a)(x - a)\] does this sound closely right.